# Universal Omni-Language Models

### v4

@@ Abstract
Recent advancements in multimodal large language models (MLLMs) have focused on integrating multiple modalities, yet their ability to simultaneously process and reason across different inputs remains underexplored. We introduce OmniBench, a novel benchmark designed to evaluate models’ ability to recognize, interpret, and reason across visual, acoustic, and textual inputs simultaneously. We define language models capable of such tri-modal processing as omni-language models (OLMs). OmniBench features high-quality human annotations that require integrated understanding across all modalities. Our evaluation reveals that: i) open-source OLMs show significant limitations in instruction-following and reasoning in tri-modal contexts; and ii) most baseline models perform poorly (around 50% accuracy) even with textual alternatives to image/audio inputs. To address these limitations, we develop OmniInstruct, an 96K-sample instruction tuning dataset for training OLMs. We advocate for developing more robust tri-modal integration techniques and training strategies to enhance OLM performance. Codes and data could be found at our repo.
Machine Learning, ICML
@@ 1 Introduction
The rapid advancement of artificial intelligence has ushered in a new era of multimodal large language models (MLLMs), capable of processing and interpreting diverse data types mainly involving images, audio, and text (Li & Lu, 2024). These models aim to emulate human-like understanding of the world by integrating information across multiple sensory modalities and learning a comprehensive context from the environment. While significant strides have been made in developing MLLMs handling two of the modalities, the ability to concurrently process and reason about the three modalities remains a frontier yet to be fully explored.
The social impact of these MLLMs is far-reaching, providing transformative capabilities for a variety of domains. In healthcare, VLMs and ALMs have contributed to diagnosing (Liu et al., 2023a; Hemdan et al., 2023), and potentially combining three modalities (Meskó, 2023). The integration of all vision, audio and text modalities is expected to significantly improve diagnostic accuracy and patient interaction, making healthcare more accessible and efficient. In urban environments, ALM can contribute to improving safety and traffic management by incorporating urban sound event detection during autonomous driving, such as recognizing audio of emergency vehicles and recognize their types or location with supplementary visual modality (Sun et al., 2021). In addition, audio contributes to biodiversity monitoring (Terenzi et al., 2021; Liang et al., 2024a) and can be greatly enhanced by MLLM’s ability to analyse both audio and video from a variety of sensors. Finally, it may help robotics or LLM agents to provide better human-computer/robotic interaction (HCI/HRI) service in day-to-day life (Liang et al., 2024b; Su et al., 2023).
@@
To address this critical need, we introduce OmniBench, a pioneering universal multimodal benchmark designed to rigorously evaluate MLLMs’ capability to recognize, interpret, and reason across visual, acoustic, and textual inputs simultaneously, which we define as the omni-understanding and reasoning ability of the omni-language models (OLMs) (Sun et al., 2024; Zhan et al., 2024; Lu et al., 2024b). For instance, one can only derive the correct answer of the sample question in Figure 1 by: 1)))) recognizing elements from the given image and audio to reconstruct the context; 2)))) interpreting the semantics and relationships among the multimodal objects according to the textual instruction formed as question and options; 3)))) reasoning and then answering with the complementary information from all the modalities. We distinguishes OmniBench by enforcing a unique constraint: accurate responses necessitate an integrated understanding and reasoning of all multimodal contexts. This approach ensures a more realistic and challenging assessment of multimodal large language models, mirroring the complex, interconnected nature of human cognition. To ensure the evaluation reliability, the development of OmniBench relies on high-quality human annotations. Furthermore, OmniBench additionally includes the answer rationales provided by the annotators to enhance the validity and ensure the benchmark aligned with human-level understanding.
@@
  • Compared to models, the results of human evaluator show not only better overall performances but a distinct distribution (e.g., much better on ”Sound Event” audios and ”Abstract Concept” tasks).
@@
2 Related Work
Multimodal Large Language Models.
Multimodal Understanding Benchmark.
Audio-Visual Understanding Datasets.
@@ 3.1 Benchmark Design
Building on the foundation of existing multimodal benchmarks, our OmniBench introduces a refined taxonomy for task categorization that effectively captures a wide range of cognitive and reasoning abilities. Our framework organizes tasks into three primary categories: (1) (temporal)-spatial entity, which includes Object Identification for recognizing distinct entities and Contextual & Environmental for discerning the setting or backdrop of the events; (2) causal inference, comprised of Story Cause Description to infer narrative drivers, Current Action & Activity to understand ongoing dynamics, and Future Plot and Purpose Inference to anticipate subsequent developments; and (3) abstract concept, involving Identity & Relationship to identify and relate entities, Text & Symbols for symbolic interpretation, and Count & Quantity for numerical reasoning. This taxonomy is designed to evaluate both foundational perceptual skills and complex cognitive processes, thereby providing a comprehensive assessment of multimodal language models’ (MLLMs) abilities to integrate and interpret diverse information sources. OmniBench includes 1142 question-answer pairs, with details on task types, text length, and the characteristics of images and audio presented in Table 1. The audio content of the dataset is categorized into speech, sound events, and music, enriching the diversity of stimuli for evaluating the models’ tri-modal capabilities and aiding in the development of future omni-language models.
@@ 3.2 Annotation Protocol
We implemented strict measures to maintain diversity across the dataset. This includes varying image and audio sources, limiting the frequency of individual speakers in audio clips to no more than five occurrences, and restricting the replication of similar instructions or questions. For instance, questions about specific environmental contexts were limited up to three samples. Importantly, annotators were required to provide rationales for correct answers, detailing the specific information that should be derived from the image and audio modalities respectively. This approach not only aided in quality inspection but also laid the groundwork for future fine-grained evaluation.
@@
Our quality control process was two-fold, including human inspection round and automatic inspection round assisted by MLLM. First, all annotated instruction-response pairs undergo cross-inspection by human annotators. Inspectors provide detailed reasons for any samples that failed to meet our stringent criteria, allowing for targeted revisions. Samples that pass human inspection are then subjected to a secondary check using a vision-language model LLaVA-1.6-34B (Liu et al., 2024a), where the the automatic quality inspection model is selected by considering trade-off between efficiency and performance. This automated process evaluates each sample under various ablation settings: image and text only, audio transcript and text only, and text only (repeated three times). Samples are only accepted if the model either rejected the task or made mistakes under these limited-information scenarios, confirming the necessity of both visual and auditory information for correct responses. We plot the distribution of the inspection frequency of the passed samples in Figure 3, where we could find that $76\%$ (868) of the passed samples do not require further modification under a well-defined annotation framework and $21.1\%$ of them requiring 1-2 times of revision. During the iterative quality checking, $9.58\%$ ($121$) QA pairs are defined as “hard to recycle” and dumped after revisions and discussions.
@@ 4 Experiment Settings
We select three groups of MLLM baselines according to the modalities available: (i) omni-language models: MIO-Instruct (Wang et al., 2024b), AnyGPT (Zhan et al., 2024), Video-SALMONN (Sun et al., 2024), UnifiedIO2 series (Lu et al., 2024b), the VITA series (Fu et al., 2024, 2025), OpenOmni (Luo et al., 2025), Baichuan-Omni-1.5 (Li et al., 2025), Qwen-2.5-Omni (Xu et al., 2025); (ii) vision-language models: InternVL-2 series (Chen et al., 2024b), Qwen2-VL series (Wang et al., 2024a), Deepseek-VL (Lu et al., 2024a), LLaVA-One-Vision series (Li et al., 2024), Cambrian series (Tong et al., 2024), Xcomposer2-4KHD (Dong et al., 2024), Idefics2 (Laurençon et al., 2024) as well as the derived Mantis-Idefics2 (Jiang et al., 2024); (iii) audio-language models: LTU series (Gong et al., 2023), Mu-LLaMA (Liu et al., 2023b), MusiLingo (Deng et al., 2023), Qwen-Audio series (Chu et al., 2023a), SALMONN-Audio (Sun et al., 2024) and Audio-Flamingo (Kong et al., 2024). We also include the API calls from proprietary models that could support image-text or audio-text inputs, including GPT4-o, Gemini Pro, Reka and Claude-3.5-Sonnet (Achiam et al., 2023; Team et al., 2023; Ormazabal et al., 2024; Anthropic, 2024). We do not conclude them as in the group of VLMs, ALMs or OLMs (even not a single model) in our context at the moment since the mechanisms behind these models are not revealed [^2] The authors conclude from an investigation on September 22, 2024.. Besides, we invite three musician with higher education background to test on the benchmark and use the average accuracy as a human expert baseline.
@@
Table 4: OLM Baselines Overall Results Grouped by Audio Type.
@@
5 Findings
We evaluate the selected baseline systems in multiple meticulously designed settings, and provide insights on the development of OLMs based on the results. All the claims stated in this section are supported by statistic significance verification, for more information please refer to Appendix C.
5.1 Results on Omni-Language Models
Overall
Table 3 demonstrates that most open-source OLM baselines surpass random guessing accuracy across various setting, and the latest developed one (Qwen-2.5-Omni) could even outperform the proprietary models. Notably, the UnifiedIO2 series demonstrates inconsistent performance scaling with model size, indicating challenges in effectively leveraging increased capacity for multimodal understanding. In contrast, Gemini-1.5-Pro and Reka-core-20240501, the two available proprietary models evaluated in this tri-modal setting, shows more promising results. Regarding the scores across audio types, the Gemini-1.5-Pro shows a more balanced performances while Reka-core-20240501 showing a lag on modeling the sound events. Moreover, the comparison of Gemini-1.5-Pro’s performance across full input context and ablated settings (image-removed and audio-removed) suggests that it effectively leverages information from all modalities to enhance its reasoning capabilities. While it demonstrates superior overall performance and balanced accuracy across audio types compared to most of the open-source alternatives, its accuracy remains below 50%.
Breakdown Results.
We present the breakdown of the performance of open-source omni-language model baselines across different audio types and task categories in the OmniBench evaluation in Table 4 and Table 5. The results reveal inconsistent performance patterns across audio types. For instance, despite overall poor performance, open-source baselines generally exhibit higher accuracy on speech audio, indicating a potential bias towards speech data. Besides, Video-Salmonn and Gemini-1.5-Pro provide better results on music subsets compared to their performance on speech and music, potentially due to their large corpus of music videos, though the music ethics of training foundation models are still under-discussion (Ma et al., 2024). Across task categories, many of the models, such as Gemini-1.5-Pro, tend to perform better on object identification and description tasks while struggling with more reasoning tasks such as plot inference and story description. This might be because visual entity recognition task is an essential component for image captioning and other type of pre-training dataset. Furthermore, some models like Gemini-1.5-Pro, Reka-core-20240501, and Video-SALMONN perform significantly badly on quantity & counting tasks compared with the rest of the tasks. But scaling up of UnifiedIO model contributes a lot to this type of task.
@@ 5.1 Results on Omni-Language Models
We invite three human annotators to evaluate on the OmniBench questions, deriving a much higher accuracy $63.19\%$ compared to all the OLMs333 Human performance is limited by time constraints, annotators’ non-native language or music knowledge gaps, and their associate-level education. . The human evaluator results hold a Fleiss’ Kappa value of $0.421$ suggests, suggesting a high level of inter-annotator agreement. Different from models, human evaluation shows a particularly good performance on ”Sound Event” among the audio types, which is reasonable since the sounds are short and straightforward without requiring complex reasoning to recognize. Moreover, human shows higher scores on ”Abstract Concept” tasks, potentially because these tasks require more reasoning.
@@ 5.2 The Effectiveness of OmniInstruct
To validate the effectiveness of OmniInstruct, we conducted experiments using MIO-instruct, a 7B parameter omni-language model previously trained only on audio-language and vision-language pairs. We evaluated the impact of supervised fine-tuning using OmniInstruct through two experimental settings. First, we created a compact subset of OmniInstruct containing 6.4K samples (approximately 7.5% of the full dataset) to enable efficient experimentation. In our first setting (MIO-Instruct-Omni-V1), we directly fine-tuned the model using its vanilla multimodal tokenizers. In the second setting (MIO-Instruct-Omni-V1-voice-filtered), we refined the approach by filtering out non-speech tokens from the audio RVQ tokenizer that only supports speech, aligning with MIO-Instruct’s original training to avoid mode collapse. The results demonstrate meaningful improvements: while the baseline MIO-Instruct achieved 24.8% accuracy on OmniBench, fine-tuning with the vanilla setting improved performance to 25.7%, and the voice-filtered approach further enhanced accuracy to 29.2%. These results indicate that even a small subset of OmniInstruct can effectively adapt existing multimodal language models to tri-modal scenarios. Notably, the superior performance of the voice-filtered approach suggests that aligning fine-tuning data distribution with a model’s distribution can yield better results.
Table 6: Results on Textual Audio Approximation Experiments. All the audios are represented in text transcript. The results are divided into groups of vision-language models and omni-models. We use the text transcript to approximate the audios in this setting. Boldface shows the best model performance, and underline shows the best open-source model.

### v6

@@ Abstract
Recent advancements in multimodal large language models (MLLMs) have aimed to integrate and interpret data across diverse modalities. However, the capacity of these models to concurrently process and reason about multiple modalities remains underexplored, partly due to the lack of comprehensive modality-wise benchmarks. We introduce OmniBench, a novel benchmark designed to rigorously evaluate models’ ability to recognize, interpret, and reason across visual, acoustic, and textual inputs simultaneously. We define language models capable of such tri-modal processing as the omni-language models (OLMs). OmniBench is distinguished by high-quality human annotations, ensuring that accurate responses require integrated understanding and reasoning across all three modalities. Our main findings reveal that: i) open-source OLMs exhibit critical limitations in instruction-following and reasoning capabilities within tri-modal contexts; and ii) most baselines models perform poorly (below 50% accuracy) even when provided with alternative textual representations of images or/and audio. These results suggest that the ability to construct a consistent context from text, image, and audio is often overlooked in existing MLLM training paradigms. To address this gap, we curate an instruction tuning dataset of 84.5K training samples, OmniInstruct, for training OLMs to adapt to tri-modal contexts. We advocate for future research to focus on developing more robust tri-modal integration techniques and training strategies to enhance OLMs. Codes, data and live leaderboard could be found at https://m-a-p.ai/OmniBench.
@@ 1 Introduction
The rapid advancement of artificial intelligence has ushered in a new era of multimodal large language models (MLLMs), capable of processing and interpreting diverse data types mainly involving images, audio, and text (Li and Lu, 2024). These models aim to emulate human-like understanding of the world by integrating information across multiple sensory modalities and learning a comprehensive context from the environment. While significant strides have been made in developing MLLMs handling two of the modalities, the ability to concurrently process and reason about the three modalities remains a frontier yet to be fully explored.
The social impact of these MLLMs is far-reaching, providing transformative capabilities for a variety of domains. In healthcare, VLMs and ALMs have contributed to diagnosing (Liu et al., 2023a; Hemdan et al., 2023), and potentially combining three modalities (Meskó, 2023). The integration of all vision, audio and text modalities is expected to significantly improve diagnostic accuracy and patient interaction, making healthcare more accessible and efficient. In urban environments, ALM can contribute to improving safety and traffic management by incorporating urban sound event detection during autonomous driving, such as recognizing audio of emergency vehicles and recognize their types or location with supplementary visual modality (Sun et al., 2021). In addition, audio contributes to biodiversity monitoring (Terenzi et al., 2021; Liang et al., 2024a) and can be greatly enhanced by MLLM’s ability to analyse both audio and video from a variety of sensors.
Finally, it may help robotics or LLM agents to provide better human-computer/robotic interaction (HCI/HRI) service in day-to-day life (Liang et al., 2024b; Su et al., 2023).
@@
To address this critical need, we introduce OmniBench, a pioneering universal multimodal benchmark designed to rigorously evaluate MLLMs’ capability to recognize, interpret, and reason across visual, acoustic, and textual inputs simultaneously [^1] Note that domain-specific modalities from field like medical, biological or financial could also be included., which we define as the omni-understanding and reasoning ability of the omni-language models (OLMs) (Sun et al., 2024; Zhan et al., 2024; Lu et al., 2024b). For instance, one can only derive the correct answer of the sample question in Figure 1 by: 1)) recognizing elements from the given image and audio to reconstruct the context; 2)) interpreting the semantics and relationships among the multimodal objects according to the textual instruction formed as question and options; 3)) reasoning and then answering with the complementary information from all the modalities. We distinguishes OmniBench by enforcing a unique constraint: accurate responses necessitate an integrated understanding and reasoning of all multimodal contexts. This approach ensures a more realistic and challenging assessment of multimodal large language models, mirroring the complex, interconnected nature of human cognition. To ensure the evaluation reliability, the development of OmniBench relies on high-quality human annotations. Furthermore, OmniBench additionally includes the answer rationales provided by the annotators to enhance the validity and ensure the benchmark aligned with human-level understanding.
@@
  • Compared to models, the results of human evaluator show not only better overall performances but a distinct distribution (e.g., much better on "Sound Event" audios and "Abstract Concept" tasks).
@@
2 Related Work
Multimodal Large Language Models.
Multimodal Understanding Benchmark.
Audio-Visual Understanding Datasets.
@@ 3.1 Benchmark Design
Building on the foundation of existing multimodal benchmarks, our OmniBench introduces a refined taxonomy for task categorization that effectively captures a wide range of cognitive and reasoning abilities. Our framework organizes tasks into three primary categories: (1) (temporal)-spatial entity, which includes Object Identification for recognizing distinct entities and Contextual & Environmental for discerning the setting or backdrop of the events; (2) causal inference, comprised of Story Cause Description to infer narrative drivers, Current Action & Activity to understand ongoing dynamics, and Future Plot and Purpose Inference to anticipate subsequent developments; and (3) abstract concept, involving Identity & Relationship to identify and relate entities, Text & Symbols for symbolic interpretation, and Count & Quantity for numerical reasoning. This taxonomy is designed to evaluate both foundational perceptual skills and complex cognitive processes, thereby providing a comprehensive assessment of multimodal language models’ (MLLMs) abilities to integrate and interpret diverse information sources. OmniBench includes 1142 question-answer pairs, with details on task types, text length, and the characteristics of images and audio presented in Table 1. It includes reasoning tasks involving: symbolic logic (e.g., matching audio cues to text clues), causal inference (e.g., predicting consequence given sound and scene), quantitative reasoning (e.g., counting musical or visual patterns), and commonsense reasoning (e.g., emotional mismatch or scene contradiction). We also highlight instances of spatial-imagery reasoning (e.g., interpreting spatial layout from soundstage cues). While our current benchmark avoids explicit chain-of-thought traces, we believe it probes deep representational understanding across modalities. Though the benchmark contains 1,142 questions, they are intentionally diverse, and we argue the inter-sample heterogeneity justifies its use for statistical comparison. Appendix D outlines our bootstrapped hypothesis testing framework, which supports generalizable model comparisons within task types. The audio content of the dataset is categorized into speech, sound events, and music, enriching the diversity of stimuli for evaluating the models’ tri-modal capabilities and aiding in the development of future omni-language models.
@@ 3.2 Annotation Protocol
We implemented strict measures to maintain diversity across the dataset. This includes varying image and audio sources, limiting the frequency of individual speakers in audio clips to no more than five occurrences, and restricting the replication of similar instructions or questions. For instance, questions about specific environmental contexts were limited up to three samples. Importantly, annotators were required to provide audio transcripts (or music captions) and rationales for correct answers, detailing the specific information that should be derived from the image and audio modalities respectively. This approach not only aided in quality inspection but also laid the groundwork for future fine-grained evaluation.
@@
Our quality control process was two-fold, including human inspection round and automatic inspection round assisted by MLLM. First, all annotated instruction-response pairs undergo cross-inspection by human annotators. Inspectors provide detailed reasons for any samples that failed to meet our stringent criteria, allowing for targeted revisions. Samples that pass human inspection are then subjected to a secondary check using a vision-language model LLaVA-1.6-34B (Liu et al., 2024a), where the the automatic quality inspection model is selected by considering trade-off between efficiency and performance. This automated process evaluates each sample under various ablation settings: image and text only, audio transcript and text only, and text only (repeated three times). Samples are only accepted if the model either rejected the task or made mistakes under these limited-information scenarios, confirming the necessity of both visual and auditory information for correct responses. We plot the distribution of the inspection frequency of the passed samples in Figure 3, where we could find that $76\%$ (868) of the passed samples do not require further modification under a well-defined annotation framework and $21.1\%$ of them requiring 1-2 times of revision. During iterative quality checking, $9.58\%$ (121121) QA pairs are defined as “hard to recycle” and dumped.
@@ 4 Experiment Settings
We select three groups of MLLM baselines according to the modalities available: (i) omni-language models: MIO-Instruct (Wang et al., 2024b), AnyGPT (Zhan et al., 2024), Video-SALMONN (Sun et al., 2024), UnifiedIO2 series (Lu et al., 2024b), the VITA series (Fu et al., 2024, 2025), OpenOmni Luo et al. (2025), Baichuan-Omni-1.5 Li et al. (2025), Qwen-2.5-Omni Xu et al. (2025); (ii) vision-language models: InternVL-2 series (Chen et al., 2024b), Qwen2-VL series (Wang et al., 2024a), Deepseek-VL (Lu et al., 2024a), LLaVA-One-Vision series (Li et al., 2024), Cambrian series (Tong et al., 2024), Xcomposer2-4KHD (Dong et al., 2024), Idefics2 (Laurençon et al., 2024) as well as the derived Mantis-Idefics2 (Jiang et al., 2024); (iii) audio-language models: LTU series (Gong et al., 2023), Mu-LLaMA (Liu et al., 2023b), MusiLingo (Deng et al., 2023), Qwen-Audio series (Chu et al., 2023a), SALMONN-Audio (Sun et al., 2024) and Audio-Flamingo (Kong et al., 2024). We also include the API calls from proprietary models that could support image-text or audio-text inputs, including GPT4-o, Gemini Pro, Reka and Claude-3.5-Sonnet (Achiam et al., 2023; Team et al., 2023; Ormazabal et al., 2024; Anthropic, 2024). We do not conclude them as in the group of VLMs, ALMs or OLMs (even not a single model) in our context at the moment since the mechanisms behind these models are not revealed  [^3] The authors conclude from an investigation on September 22, 2024.. Besides, we invite three musician with higher education background to test on the benchmark and use the average accuracy as a human expert baseline.
@@
Human Expert 74.03% - -
5 Findings
We evaluate the selected baseline systems in multiple meticulously designed settings, and provide insights on the development of OLMs based on the results. All the claims stated in this section are supported by statistic significance verification, for more information please refer to Appendix D.
Figure 4: OLM Baselines Overall Results Grouped by Audio Type.
@@ 5 Findings
5.1 Results on Omni-Language Models
Overall
Table 3 demonstrates that most open-source OLM baselines surpass random guessing accuracy across various setting, and the latest developed one (Qwen-2.5-Omni) could even outperform the proprietary models. Notably, the UnifiedIO2 series demonstrates inconsistent performance scaling with model size, indicating challenges in effectively leveraging increased capacity for multimodal understanding. In contrast, Gemini-1.5-Pro and Reka-core-20240501, the two available proprietary models evaluated in this tri-modal setting, shows more promising results. Regarding the scores across audio types, the Gemini-1.5-Pro shows a more balanced performances while Reka-core-20240501 showing a lag on modeling the sound events. Moreover, the comparison of Gemini-1.5-Pro’s performance across full input context and ablated settings (image-removed and audio-removed) suggests that it effectively leverages information from all modalities to enhance its reasoning capabilities. While it demonstrates superior overall performance and balanced accuracy across audio types compared to most of the open-source alternatives, its accuracy remains below 50%. OmniBench’s best-performing model (Qwen-Omni) also leads in audio reasoning. Interestingly, it combines separate audio and vision instruction tuning, suggesting that fusing different bimodal corpora can generalize to tri-modal settings. Although most top-performing models still rely on modality-specific encoders, emerging models like AnyGPT—which use a universal tokenizer for multiple modalities—represent a more integrated architecture. We hypothesize that GPT-4o may follow this direction, and believe such designs will be necessary to scale OLMs to more complex understanding and generation tasks.
Breakdown Results.
We present the breakdown of the performance of open-source omni-language model baselines across different audio types and task categories in the OmniBench evaluation in Figure 4 and Table 4. The results reveal inconsistent performance patterns across audio types. For instance, despite overall poor performance, open-source baselines generally exhibit higher accuracy on speech audio, indicating a potential bias towards speech data. Besides, Video-Salmonn and Gemini-1.5-Pro provide better results on music subsets compared to their performance on speech and music, potentially due to their large corpus of music videos, though the music ethics of training foundation models are still under-discussion (Ma et al., 2024).
@@ 5.1 Results on Omni-Language Models
Across task categories, many of the models, such as Gemini-1.5-Pro, tend to perform better on object identification and description tasks while struggling with more reasoning tasks such as plot inference and story description. This might be because visual entity recognition task is an essential component for image captioning and other type of pre-training dataset. Furthermore, some models like Gemini-1.5-Pro, Reka-core-20240501, and Video-SALMONN perform significantly badly on quantity & counting tasks compared with the rest of the tasks. But scaling up of UnifiedIO model contributes a lot to this type of task.
Results on Music-related Questions
Model underperform on music subset in Figure 4 and Table 5 is notable and multifactorial. While OmniBench is carefully curated to balance genre and linguistic diversity (e.g., avoiding overrepresentation of English pop), many audio-capable LLMs suffer from training-data bias. For instance, Qwen2-Audio leverages MTG-Jamendo for music tagging, yielding strong results there but poor generalization to other datasets like MTT (Ma et al., 2025). In contrast to speech, music is more restricted by copyright and expensive to annotate for supervised tasks.
@@
We invite three human annotators to evaluate on the OmniBench questions, deriving a much higher accuracy $63.19\%$ compared to all the OLMs [^4] Human performance is limited by time constraints, annotators’ non-native language or music knowledge gaps, and their associate-level education. . The human evaluator results hold a Fleiss’ Kappa value of $0.421$ suggests, suggesting a high level of inter-annotator agreement. Different from models, human evaluation shows a particularly good performance on "Sound Event" among the audio types, which is reasonable since the sounds are short and straightforward without requiring complex reasoning to recognize. Moreover, human shows higher scores on "Abstract Concept" tasks, potentially because these tasks require more reasoning.
@@ 5.2 The Effectiveness of OmniInstruct
To validate the effectiveness of OmniInstruct, we conducted experiments using MIO-instruct, a 7B parameter omni-language model previously trained only on audio-language and vision-language pairs. We evaluated the impact of supervised fine-tuning using OmniInstruct through two experimental settings. First, we created a compact subset of OmniInstruct containing 6.4K samples (approximately 7.5% of the full dataset) to enable efficient experimentation. In our first setting (MIO-Instruct-Omni-V1), we directly fine-tuned the model using its vanilla multimodal tokenizers. In the second setting (MIO-Instruct-Omni-V1-voice-filtered), we refined the approach by filtering out non-speech tokens from the audio RVQ tokenizer that only supports speech, aligning with MIO-Instruct’s original training to avoid mode collapse. The results demonstrate meaningful improvements: while the baseline MIO-Instruct achieved 24.8% accuracy on OmniBench, fine-tuning with the vanilla setting improved performance to 25.7%, and the voice-filtered approach further enhanced accuracy to 29.2%. When we leverage the full OminiInstruct to LoRAfine-tune MiniCPM-o-2.6, the model improve its performance from 40.5% on to 45.9% on OmniBench. These results indicate that even a small subset of OmniInstruct can effectively adapt existing multimodal language models to tri-modal scenarios. Notably, the superior performance of the voice-filtered approach suggests that aligning fine-tuning data distribution with a model’s distribution can yield better results.
Figure 5: Results on Textual Audio Approximation Experiments. All the audios are represented in text transcript. The results are divided into groups of vision-language models and omni-models. We use the text transcript to approximate the audios in this setting. Boldface shows the best model performance, and underline shows the best open-source model.
@@
5.3 Textual Approximation on Images and Audios
As the absence of strong OLM baselines, we further introduce the text alternatives of images ($I^{\prime}$) and audios ($A^{\prime}$) to embrace more dual-modal MLLMs to analyze current research progress.
Performance Changes of Open OLMs.
Performances of Dual-modal MLLMs.
In the setting of using text as the alternatives of audios and images, the VLMs show generally better results than ALMs (Figure 5 vs Table 5) even compared with open-source model with similar model size [^5] The results of pure text $(I^{\prime},A^{\prime})$ setting are placed at Table 6, Appendix A.. This could be caused by : 1) more available research resources have been put in VLMs to develop datasets and cross-modality alignment architectures, leading to higher instruction following rate and accuracy compared to ALMs; 2) the audio data are naturally harder (and hence more expensive) to annotate; and 3) audio typically has longer sequence tokens and requires more computational resource compared to text and image, making it harder to scale up. If $I^{\prime}$ and $A^{\prime}$ have the information loss ratio when converted from $I$ and $A$, it seems to be easier for the researchers to train the future omni-language models from exisiting VLMs rather than ALMs. Besides, we can observe Claude-3.5 and GPT-4o are generally the best two VLMs, significantly better compared to open-source VLMs. Qwen2- audio and Gemini are the two best ALMs in speech, Qwen2- audio is the best in sound, and Audio-SALMONN is music. Moreover, we can see significant differences in different types of audio, i.e., LTU and audio-flamingo are worse for music compared to speech, while Qwen-audio, which includes music on pre-training, provides better results on music compared to speech. And MusiLingo only uses music for pre-training performs worse in speech and audio.